Let: A = no. of letters of your given name

B = no. of letters of your middle name


A) Find the speed of a wave if its if its frequency is A Hz and its wavelength is B meters.

B) An radio transmitter produces a wave that has a wavelength of 600 m and travels at 200 m/s. What is its frequency?

C) Research on the cause why the sky is blue and sunsets mostly red, orange and yellow.


1
Expert's answer
2021-05-27T18:52:02-0400

(a).According to the question frequency"=A"and Wavelength"=B"

"\Rightarrow Speed=Frequency\times Wavelength"

"\Rightarrow Speed=A\times B"

Speed of Wave will be "AB"


(b).We know that "c=200m/s,\lambda=600m"

"\Rightarrow \nu=\dfrac{c}{\lambda}\\"


"\Rightarrow \nu= \dfrac{200}{600}=0.333s^{-1}"

Frequency of radio transmitter will be "0.333s^{-1}"


(c).Small particles of dust and pollution in the air can contribute to (and sometimes even enhance) these colors, but the primary cause of a blue sky and orange/red sunsets or sunrises is scattering by the gas molecules that make up our atmosphere. In this way, clouds act as a screen on which nature's colors are painted.